OpenStudy (anonymous):

The strength of an electric field at 2.4 m from a point charge is 250.0 N/C. What is the magnitude of the source charge? **How can I show the work for this? Thank you :)

OpenStudy (michele_laino):

here we have to apply this formula: \[E = K\frac{Q}{{{r^2}}}\] where E is the electric field, Q is the electric charge, and r is the distance from the electric charge to the point of observation

OpenStudy (michele_laino):

as usually K = 9*10^9 N*m^2/Coulomb^2

OpenStudy (michele_laino):

so we have to write the inverse formula, to find Q. What is that formula?

OpenStudy (anonymous):

ohh okay! and i'm not sure :( i forgot :(

OpenStudy (michele_laino):

here is it: \[\Large Q = \frac{{E{r^2}}}{K}\]

OpenStudy (michele_laino):

so substituting your data we get: \[\Large Q = \frac{{E{r^2}}}{K} = \frac{{250 \times {{\left( {2.4} \right)}^2}}}{{9 \times {{10}^9}}} = ...?\]

OpenStudy (anonymous):

ahh okay, that's right!! :) so then 1440/9*10^9 ? i'm not sure what that gets? My calculator says 1.6 E^-7.. is that correct?

OpenStudy (michele_laino):

ok! we have: \[\large Q = \frac{{E{r^2}}}{K} = \frac{{250 \times {{\left( {2.4} \right)}^2}}}{{9 \times {{10}^9}}} = 1.6 \times {10^{ - 7}} = 0.16 \times {10^{ - 6}} = 0.16\mu C\]

OpenStudy (anonymous):

ohhh so you condensed it! i see:) yay!! so what happens next? is there more? :O or is 0.16µC the final answer?

OpenStudy (michele_laino):

It is the final answer!

OpenStudy (anonymous):

yay!! thank you:)

OpenStudy (michele_laino):

thank you! :)